# Disentangling modified gravity from a dark force with gravitational redshift

### v1

@@ Abstract
The standard approach to test for deviations from General Relativity on cosmological scales is to combine measurements of the growth rate of structure with gravitational lensing. In this study, we show that this method suffers from an important limitation with regard to these two probes: models of dark matter with additional interactions can lead to the very same observational signatures found in modified gravity (and viceversa). Using synthetic data of redshift-space distortions, weak lensing, and cosmic microwave background, we demonstrate that this degeneracy is inevitable between modifications of gravity and a dark fifth force. We then show that the coming generation of surveys, in particular the Square Kilometer Array, will allow us to break the degeneracy between such models through measurements of gravitational redshift. Performing a Markov Chain Monte Carlo analysis of the synthetic data set, we quantify the extent to which gravitational redshift can distinguish between two representative classes of models, Generalized Brans-Dicke (modified gravity) and Coupled Quintessence (fifth force).
@@ I Introduction
In addition to dark energy, several observations have strongly suggested the presence of another obscure component, dark matter, based on its gravitational influence on visible matter and light. In the standard cosmological model, dark matter takes the form of a slow-moving (“cold”), non-interacting, non-baryonic species. Since such a particle has not been detected yet, several extensions to the simple CDM paradigm have been proposed. These include scenarios where dark matter interacts weakly with particles of the Standard Model Barkana (2018); Schewtschenko et al. (2016), with the dark sector (e.g. dark radiation) Diamanti et al. (2013); Buen-Abad et al. (2015), or through self-interactions via a fifth (non-gravitational) force Spergel and Steinhardt (2000); Pettorino et al. (2012); Costa et al. (2015); Archidiacono et al. (2022).
@@
To test for gravity modifications and dark matter interactions, the standard approach is to employ measurements of the growth rate of cosmic structure obtained through redshift-space distortions (RSD) (see e.g. Refs. Blake et al. (2011); Howlett et al. (2015); Alam et al. (2017, 2021)). Since gravity governs the way cosmic structure evolves with time, deviations from GR generically modify the growth rate. Similarly, the growth rate would also be affected by any new force acting on the dark matter particles, as this would have an impact on structure formation.
@@
A modification of gravity would involve a deviation in the Poisson equation affecting the evolution of all cosmic components, whereas a fifth force would only modify the Euler equation for dark matter. Current analyses are testing these two scenarios separately. The community working on gravity modifications usually assumes that dark matter is a cold non-interacting particle and uses the growth rate of structure to test theories beyond GR, see e.g. Ref. Abbott et al. (2023). By contrast, the community working on non-standard dark matter models typically assumes the validity of GR and uses the growth rate to constrain the strength of dark matter interactions Pourtsidou et al. (2013); Buen-Abad et al. (2018); Bottaro et al. (2023). As we do not know which scenario (if any) is correct, it is important to ask whether it is possible to test for both kinds of modifications at the same time and disentangle them from one another.
@@
In Ref. Bonvin and Pogosian (2023), it was shown that this cannot be achieved with current observables. The growth rate of structure is affected by the modifications in both scenarios, potentially in the very same way. On the other hand, weak lensing, which is sensitive to the sum of the two gravitational potentials describing the geometry of the Universe, is also generically unable to discriminate between the two cases. The degeneracy can however be broken by considering measurements of an observable accessible by the coming generation of galaxy surveys: gravitational redshift Sobral-Blanco and Bonvin (2021, 2022). This effect, originally predicted by Einstein Einstein (1916), directly probes the gravitational potential encoding the distortion of time and entering the Euler equation. This yields an immediate test of modified gravity and provides a way to distinguish it from the presence of a fifth force acting on dark matter.
@@
In this work, we quantify the capability of gravitational redshift to break the degeneracy between gravity modifications and a dark fifth force, performing a Markov Chain Monte Carlo (MCMC) analysis on synthetic data, based on the survey specifications of the Square Kilometer Array Phase 2 (SKA2) Bull (2016). We consider two specific models: the Symmetron modified gravity model Hinterbichler and Khoury (2010), which is a scalar-tensor theory of Generalized Brans-Dicke (GBD) type, and a Coupled Quintessence (CQ) model Amendola et al. (2018); Barros et al. (2019) with exactly the same form of the coupling and the potential as the Symmetron, but with only dark matter coupled to the scalar field. We show that the parameters of the two models are fully degenerate in an analysis including RSD, weak lensing and Cosmic Microwave Background (CMB) data. We then demonstrate that including gravitational redshift into the analysis breaks the degeneracy, and we determine how large the deviations from $\Lambda$CDM need to be for this observable to be effective.
@@
The rest of this paper is organized as follows. In Sec. II, we present the GBD and CQ models considered in our analysis, highlighting their degeneracies in RSD and weak lensing data. We then describe the galaxy clustering, weak lensing and CMB observables included in the analysis in Sec. III, and discuss the details of the numerical investigation in Sec. IV. Finally, we present the results in Sec. V and conclude in Sec. VI. We include some additional details on the GBD and CQ models in Appendix A and on the specifications for the galaxy clustering observables in Appendix B.
II The two degenerate scenarios
@@ II.1 GBD and CQ
In order to illustrate the disentangling power of gravitational redshift, we compare a modified gravity model with a scenario involving a dark fifth force acting on dark matter. The argument that follows is fully general and applicable to any model belonging to these categories, but for the sake of concreteness, we focus on a GBD scalar-tensor theory and a CQ model, following Ref. Bonvin and Pogosian (2023). Both scenarios involve an additional scalar field, according to the actions given in Appendix A. In the GBD case, the scalar field has the same conformal coupling to all cosmic components, whereas in CQ the coupling only involves dark matter (in the form of CDM) and propagates a dark force of non-gravitational origin. We denote the coupling strengths with $\beta_{1}$ and $\beta_{2}$ and the scalar field masses with $m_{1}$ and $m_{2}$ in the GBD and CQ models, respectively.
@@ II.2 Impact on the growth of cosmic structure
where $G$ is the Newton constant and we have defined $\tilde{\beta}^{2}_{i}=\beta^{2}_{i}/(8\pi G)$, for $i=1,2$. We remark that the only difference between the two expressions is the term $\left(\rho_{c}/\rho\right)^{2}\left(\delta_{c}/\delta\right)$ suppressing the value of $G_{\rm eff}$ in the CQ case, due to the fact that the coupling only affects CDM in this scenario. However, such a difference can be absorbed into the unknown value of the coupling $\beta_{2}$. Therefore, measurements of the growth of cosmic structure that only constrain $G_{\rm eff}$ cannot distinguish between a scenario where $\beta_{1}=0$ (GR is valid) and $\beta_{2}\neq 0$ (CDM experiences a fifth force), and a scenario where $\beta_{1}\neq 0$ (gravity is modified) and $\beta_{2}=0$ (there is no dark fifth force). In general, if we analyze the data allowing for both $\beta_{1}$ and $\beta_{2}$ to vary, we see from Eqs. (3) and (4) that one can only constrain the combination $\beta_{1}^{2}+\left(\rho_{c}/\rho\right)^{2}\left(\delta_{c}/\delta\right)\beta_{2}^{2}$. While the shape of the degeneracy depends on the specific models considered, deviations in the Poisson and the Euler equations are generically indistinguishable through measurements of the growth of structure Castello et al. (2022).
@@ II.3 Impact on weak lensing
The other key large-scale structure observable is weak lensing, measured in cosmic shear, magnification of high-redshift galaxies, and the lensing of the CMB. Weak lensing is directly sensitive to the Weyl potential, i.e. the sum of the two gravitational potentials $\Phi$ and $\Psi$ Dodelson (2003). However, as shown in Ref. Bonvin and Pogosian (2023), measurements of the Weyl potential cannot disentangle GBD from CQ, as in both scenarios we have the following Poisson-like relation [^1] This assumes that $A^{2}\approx 1$ in the GBD action in Eq. (A), which is required for the screening mechanism to be effective Joyce et al. (2015). As discussed in Appendix A, this assumption has no impact on the arguments presented.,
@@ II.4 The setup for the analysis
for $a>a_{\star}$. We fix $a_{\star}=0.5$ for both scenarios and study the constraints on the parameters $\beta_{*1}$ and $\beta_{*2}$ for different values of the masses that set the Compton wavelength of the scalar field.
@@ III Observables
We consider three observables to constrain the GBD and CQ models: galaxy clustering, weak lensing and the CMB.
@@ III.1 Galaxy clustering
where $N$ is the number of galaxies in a pixel centered in direction $\mathbf{\hat{n}}$ and at redshift $z$, and $\bar{N}$ denotes the average number of galaxies inside the pixel. The observable $\Delta$ can be expressed within linear perturbation theory as Bonvin and Durrer (2011); Challinor and Lewis (2011); Yoo et al. (2009)
@@
where $r$ denotes the comoving distance to the galaxies, $b$ is the galaxy bias, $s$ is the magnification bias and $f^{\rm evol}$ is the evolution bias. The dominant contribution to $\Delta$ arises from the first two terms, which respectively encode the effect of matter density perturbations and RSD Kaiser (1987). These are the only two contributions that are measurable with current data. The other terms are relativistic corrections suppressed on sub-horizon scales by a factor of ${\mathcal{H}}/{k}$, including the Doppler terms and the gravitational redshift effect given by the radial derivative of $\Psi$ [^2] Note that $\Delta$ is also affected by other relativistic effects suppressed by $(\mathcal{H}/k)^{2}$ and by gravitational lensing, whose impact is negligible in the redshift range considered in this work Jelic-Cizmek et al. (2021); Lepori et al. (2022).
We can extract information from $\Delta(\mathbf{\hat{n}},z)$ by measuring its two-point correlation function $\xi\equiv\langle\Delta(\mathbf{\hat{n}},z)\Delta(\mathbf{\hat{n}}^{\prime},z^{\prime})\rangle$. The density and RSD terms in Eq. (9) generate three even multipoles in the correlation function: a monopole, a quadrupole and a hexadecapole. Relativistic corrections to these three multipoles have been shown to be negligible, being suppressed by $(\mathcal{H}/k)^{2}$ relative to the density and RSD contributions Jelic-Cizmek et al. (2021). One can however exploit the anti-symmetry generated by relativistic effects. This is manifested in the presence of odd multipoles, most notably a dipole, when cross-correlating two different populations of galaxies Bonvin et al. (2014); Bonvin (2014); Gaztanaga et al. (2017); Yoo et al. (2012); Croft (2013). Since the dipole is sensitive to the effect of gravitational redshift, which is suppressed by only a single power of $\mathcal{H}/k$, measuring it can help to disentangle modified gravity from a dark fifth force, as we show below.
The cross-correlation between two populations of galaxies with different luminosities—a bright and faint sample labelled B and F, respectively—generates the following even multipoles in the flat-sky approximation:
@@
Here, $j_{\ell}$ denotes the spherical Bessel function of order $\ell$ and $P_{\delta\delta},P_{\theta\theta}$ and $P_{\delta\theta}$ are the auto- and cross-power spectra, which depend on $k$ and $z$. The velocity divergence $\theta$ is defined in Fourier space by $\mathbf{V}=i(\mathbf{k}/k^{2})\,\theta$. The dipole is given by
@@
$-\frac{1}{5\pi^{2}}\int\mathrm{d}k\,k^{2}\frac{d}{r}\left(\frac{1}{\mathcal{H}}\right)(b_{\mathrm{B}}-b_{\mathrm{F}})P_{\delta\theta}(k,z)j_{2}(kd)\,,$ (13)
@@
The first term in Eq. (III.1), proportional to $P_{\delta\Psi}$, is the contribution from gravitational redshift. This key term is not present in the even multipoles and provides the information needed to disentangle the GBD and CQ scenarios. The last term in Eq. (III.1) is a wide-angle correction, which is needed for a consistent treatment of the dipole Bonvin et al. (2016). [^3] Note that wide-angle corrections also enter the even multipoles, but can be neglected since they are more suppressed compared to those in the dipole Reimberg et al. (2016); Bonvin et al. (2014). Note that anti-symmetry can only be probed by cross-correlating distinct populations, meaning that the dipole signal is nonzero only for $\mathrm{B}\neq\mathrm{F}$.

### v2

@@ Abstract
The standard approach to test for deviations from general relativity on cosmological scales is to combine measurements of the growth rate of structure with gravitational lensing. In this study, we show that this method suffers from an important limitation with regard to these two probes: models of dark matter with additional interactions can lead to the very same observational signatures found in modified gravity and vice versa. Using synthetic data of redshift-space distortions, weak lensing, and cosmic microwave background, we demonstrate that this degeneracy is inevitable between modifications of gravity and a dark fifth force. We then show that the coming generation of surveys, in particular the Square Kilometre Array, will allow us to break the degeneracy between such models through measurements of gravitational redshift. Performing a Markov chain Monte Carlo analysis of the synthetic dataset, we quantify the extent to which gravitational redshift can distinguish between two representative classes of models, generalized Brans-Dicke (modified gravity) and coupled quintessence (fifth force).
@@ I Introduction
In addition to dark energy, a range of observations have strongly suggested the presence of another obscure component, dark matter, based on its gravitational influence on visible matter and light. In the standard cosmological model, dark matter takes the form of a slow-moving (“cold”), noninteracting, nonbaryonic species. Although CDM has proved remarkably successful in explaining structure formation on large scales, its observational status on small scales is less clear Bullock and Boylan-Kolchin (2017); Salucci (2019). These issues, together with the fact that dark matter has yet to be directly detected, have motivated a plethora of alternatives and extensions to the simple CDM paradigm. For example, dark matter has been proposed to interact weakly with particles of the Standard Model Barkana (2018); Schewtschenko et al. (2016), with the dark sector (e.g. dark radiation) Diamanti et al. (2013); Buen-Abad et al. (2015), or through self-interactions via a fifth (nongravitational) force Spergel and Steinhardt (2000); Pettorino et al. (2012); Costa et al. (2015); Archidiacono et al. (2022).
@@
To test for gravity modifications and dark matter interactions, the standard approach is to employ measurements of the growth rate of cosmic structure obtained through redshift-space distortions (RSD) (see e.g. Refs. Blake et al. (2011); Howlett et al. (2015); Alam et al. (2017, 2021); Nguyen et al. (2023)). Since gravity governs the way cosmic structure evolves with time, deviations from GR generically modify the growth rate. Similarly, the growth rate would also be affected by any new force acting on the dark matter particles, as this would have an impact on structure formation.
@@
A modification of gravity would involve a deviation in the Poisson equation affecting the evolution of all cosmic components, whereas a fifth force would only modify the Euler equation for dark matter. Current analyses are testing these two scenarios separately. The community working on gravity modifications usually assumes that dark matter is a cold noninteracting particle and uses the growth rate of structure to test theories beyond GR, see e.g. Ref. Abbott et al. (2023). By contrast, the community working on nonstandard dark matter models typically assumes the validity of GR and uses the growth rate to constrain the strength of dark matter interactions Pourtsidou et al. (2013); Buen-Abad et al. (2018); Bottaro et al. (2024). As we do not know which scenario (if any) is correct, it is important to ask whether it is possible to test for both kinds of modifications at the same time and disentangle them from one another.
In Ref. Bonvin and Pogosian (2023), it was shown that this cannot be achieved with current observables. The growth rate of structure is affected by the modifications in both scenarios, potentially in the very same way. On the other hand, weak lensing, which is sensitive to the sum of the two gravitational potentials describing the geometry of the Universe, is also generically unable to discriminate between the two cases. While it is possible to look for observational signatures of specific models (see e.g. Refs. Joyce et al. (2015); Burrage et al. (2019); Käding (2023)), these two scenarios cannot be disentangled without knowing a priori which kind of modifications (if any) are present in the data.
@@
Reference Bonvin and Pogosian (2023), however, pointed out that the degeneracy can be broken by considering measurements of an observable accessible by the coming generation of galaxy surveys: gravitational redshift Sobral-Blanco and Bonvin (2021, 2022). This effect, originally predicted by Einstein Einstein (1916), directly probes the gravitational potential encoding the distortion of time and entering the Euler equation. This yields an immediate test of modified gravity and provides a way to distinguish it from the presence of a fifth force acting on dark matter.
@@
In this work, we quantify the capability of gravitational redshift to break the degeneracy between gravity modifications and a dark fifth force, performing a Markov chain Monte Carlo (MCMC) analysis on synthetic data, based on the survey specifications of the Square Kilometre Array Phase 2 (SKA2) Bull (2016). We consider two specific models: the symmetron modified gravity model Hinterbichler and Khoury (2010), which is a scalar-tensor theory of generalized Brans-Dicke (GBD) type, and a coupled quintessence (CQ) model Amendola et al. (2018); Barros et al. (2019) with exactly the same form of the coupling and the potential as the symmetron, but with only dark matter coupled to the scalar field. We show that the parameters of the two models are fully degenerate in an analysis including RSD, weak lensing, and cosmic microwave background (CMB) data. We then demonstrate that including gravitational redshift into the analysis breaks the degeneracy, and we determine how large the deviations from $\Lambda$CDM need to be for this observable to be effective.
@@
The rest of this paper is organized as follows. In Sec. II, we present the GBD and CQ models considered in our analysis, highlighting their degeneracies in RSD and weak lensing data. We then describe the galaxy clustering, weak lensing and CMB observables included in the analysis in Sec. III and discuss the details of the numerical investigation in Sec. IV. Finally, we present the results in Sec. V and conclude in Sec. VI. We include some additional details on the GBD and CQ models in Appendix A and on the specifications for the galaxy clustering observables in Appendix B.
II Two degenerate scenarios
@@ II.1 GBD and CQ
In order to illustrate the disentangling power of gravitational redshift, we compare a modified gravity model with a scenario involving a dark fifth force acting on dark matter. The argument that follows is fully general and applicable to any model belonging to these categories, but for the sake of concreteness, we focus on a GBD scalar-tensor theory and a CQ model, following Ref. Bonvin and Pogosian (2023). Both scenarios involve an additional scalar field, according to the actions given in Appendix A. In the GBD case, the scalar field has the same conformal coupling to all cosmic components, whereas in CQ the coupling only involves dark matter (in the form of CDM) and propagates a dark force of nongravitational origin. We denote the coupling strengths with $\beta_{1}$ and $\beta_{2}$ and the scalar field masses with $m_{1}$ and $m_{2}$ in the GBD and CQ models, respectively.
@@ II.2 Impact on the growth of cosmic structure
where $G$ is the Newton constant and we have defined $\tilde{\beta}^{2}_{i}=\beta^{2}_{i}/(8\pi G)$, for $i=1,2$. We remark that the only difference between the two expressions is the term $\left(\rho_{c}/\rho\right)^{2}\left(\delta_{c}/\delta\right)$ suppressing the value of $G_{\rm eff}$ in the CQ case, due to the fact that the coupling only affects CDM in this scenario. However, such a difference can be absorbed into the unknown value of the coupling $\beta_{2}$. Therefore, measurements of the growth of cosmic structure that only constrain $G_{\rm eff}$ cannot distinguish between a scenario where $\beta_{1}=0$ (GR is valid) and $\beta_{2}\neq 0$ (CDM experiences a fifth force), and a scenario where $\beta_{1}\neq 0$ (gravity is modified) and $\beta_{2}=0$ (there is no dark fifth force). In general, if we analyze the data allowing for both $\beta_{1}$ and $\beta_{2}$ to vary, we are sensitive to $G_{\rm eff}=G_{\rm eff}^{\rm GBD}+G_{\rm eff}^{\rm CQ}$. Assuming $m_{1}=m_{2}\equiv m$, this yields
$G_{\rm eff}=G\left[1+{2k^{2}\over a^{2}m^{2}+k^{2}}\left(\tilde{\beta}_{1}^{2}+\left(\rho_{c}\over\rho\right)^{2}\left(\delta_{c}\over\delta\right)\tilde{\beta}_{2}^{2}\right)\right],$ (5)
indicating that measurements of the growth of structure can only constrain the combination $\tilde{\beta}_{1}^{2}+\left(\rho_{c}/\rho\right)^{2}\left(\delta_{c}/\delta\right)\tilde{\beta_{2}}^{2}$. While the shape of the degeneracy depends on the specific models considered, deviations in the Poisson and the Euler equations are generically indistinguishable through measurements of the growth of structure Castello et al. (2022).
@@ II.3 Impact on weak lensing
The other key large-scale structure observable is weak lensing, measured in cosmic shear, magnification of high-redshift galaxies, and the lensing of the CMB. Weak lensing is directly sensitive to the Weyl potential, i.e. the sum of the two gravitational potentials $\Phi$ and $\Psi$ Dodelson (2003). However, as shown in Ref. Bonvin and Pogosian (2023), measurements of the Weyl potential cannot disentangle GBD from CQ, as in both scenarios we have the following Poisson-like relation, [^1] This assumes that $A^{2}\approx 1$ in the GBD action in Eq. (A), which is required for the screening mechanism to be effective Joyce et al. (2015). As discussed in Appendix A, this assumption has no impact on the arguments presented.
@@ II.4 Setup for the analysis
for $a>a_{\star}$. We fix $a_{\star}=0.5$ for both scenarios and study the constraints on the parameters $\beta_{*1}$ and $\beta_{*2}$ for different values of the masses that set the Compton wavelength of the scalar field. For the purpose of our demonstration, we choose a scenario where the two kinds of modifications have exactly the same time evolution, thus allowing us to treat both on an equal footing.
@@ III Observables
We consider three observables to constrain the GBD and CQ models: galaxy clustering, weak lensing, and the CMB.
@@ III.1 Galaxy clustering
where $N$ is the number of galaxies in a pixel centered in direction $\mathbf{\hat{n}}$ and at redshift $z$ and $\bar{N}$ denotes the average number of galaxies inside the pixel. The observable $\Delta$ can be expressed within linear perturbation theory as Bonvin and Durrer (2011); Challinor and Lewis (2011); Yoo et al. (2009)
@@
where $r$ denotes the comoving distance to the galaxies, $b$ is the galaxy bias, $s$ is the magnification bias and $f^{\rm evol}$ is the evolution bias. The dominant contribution to $\Delta$ arises from the first two terms, which respectively encode the effect of matter density perturbations and RSD Kaiser (1987). These are the only two contributions that are measurable with current data. The other terms are relativistic corrections suppressed on subhorizon scales by a factor of ${\mathcal{H}}/{k}$, including the Doppler terms and the gravitational redshift effect given by the radial derivative of $\Psi$.  [^2] Note that $\Delta$ is also affected by other relativistic effects suppressed by $(\mathcal{H}/k)^{2}$ and by gravitational lensing, whose impact is negligible in the redshift range considered in this work Jelic-Cizmek et al. (2021); Lepori et al. (2022).
We can extract information from $\Delta(\mathbf{\hat{n}},z)$ by measuring its two-point correlation function $\xi\equiv\langle\Delta(\mathbf{\hat{n}},z)\Delta(\mathbf{\hat{n}}^{\prime},z^{\prime})\rangle$. The density and RSD terms in Eq. (10) generate three even multipoles in the correlation function: a monopole, a quadrupole and a hexadecapole, which probe the growth of structure and are the key quantities measured in a standard RSD analysis. Relativistic corrections to these three multipoles have been shown to be negligible, being suppressed by $(\mathcal{H}/k)^{2}$ relative to the density and RSD contributions Jelic-Cizmek et al. (2021). One can however exploit the antisymmetry generated by relativistic effects. This is manifested in the presence of odd multipoles, most notably a dipole, when cross-correlating two different populations of galaxies Bonvin et al. (2014); Bonvin (2014); Gaztanaga et al. (2017); McDonald (2009); Yoo et al. (2012); Croft (2013). Since the dipole is sensitive to the effect of gravitational redshift, which is suppressed by only a single power of $\mathcal{H}/k$, measuring it can help to disentangle modified gravity from a dark fifth force, as we show below.
The cross-correlation between two populations of galaxies with different luminosities—a bright and faint sample labeled B and F, respectively—generates the following even multipoles in the flat-sky approximation:
@@
Here, $j_{\ell}$ denotes the spherical Bessel function of order $\ell$, and $P_{\delta\delta},P_{\theta\theta}$, and $P_{\delta\theta}$ are the auto- and cross-power spectra, which depend on $k$ and $z$. The velocity divergence $\theta$ is defined in Fourier space by $\mathbf{V}=i(\mathbf{k}/k^{2})\,\theta$. The dipole is given by
@@
$+\frac{1}{5\pi^{2}}\int\mathrm{d}k\,k^{2}\frac{d}{r}\left(\frac{1}{\mathcal{H}}\right)(b_{\mathrm{B}}-b_{\mathrm{F}})P_{\delta\theta}(k,z)j_{2}(kd)\,,$ (14)
@@
The first term in Eq. (III.1), proportional to $P_{\delta\Psi}$, is the contribution from gravitational redshift. This key term is not present in the even multipoles and provides the information needed to disentangle the GBD and CQ scenarios. The last term in Eq. (III.1) is a wide-angle correction, which is needed for a consistent treatment of the dipole Bonvin et al. (2016). [^3] Note that wide-angle corrections also enter the even multipoles but can be neglected since they are more suppressed compared to those in the dipole Reimberg et al. (2016); Bonvin et al. (2014). Note that antisymmetry can only be probed by cross-correlating distinct populations, meaning that the dipole signal is nonzero only for $\mathrm{B}\neq\mathrm{F}$.